OpenStudy (anonymous):

solve by substituting a variable to simplify the equation: x^4+2x^2=8 How do i get the answer; + or - the square root of 2 and + or - 2i

OpenStudy (anonymous):

Try substituting y=x^2. So your equation is now: y^2+2y=8

OpenStudy (anonymous):

okay then y^2+2y-8=0?

OpenStudy (anonymous):

Yup! You're on the right track!

OpenStudy (anonymous):

okay but where did the "i" come from in the answer?

OpenStudy (anonymous):

Well, when you factor you get: (y+4)(y-2)=0 So y=-4 or y=2 which means x^2=-4 or x^2=2 square root both sides and you get x=+-2i and x=+-sqrt2

OpenStudy (anonymous):

thanks!